A Bag Contains 5 Red Marbles and 5 blue marbles. Without Replacement what is the probability of having two successive selections of Red Followed by two Successive selections of blue? I think since 5+5=10 the fraction would be 5/10 which =1/2, and than 2+2=4 so 4/2 would be the answer?
Answer by stanbon(75887)

A Bag Contains 5 Red Marbles and 5 blue marbles. Without Replacement what is the probability of having two successive selections of Red Followed by two Successive selections of blue?
----
P(rrbb) = (5/10)(4/9)(5/8)(4/7) = 400/5040 = 0.0794
====================
Cheers,
Stan H.
====================